Find a positive number that is 20 less than its square
Answer by ankor@dixie-net.com(22740)

Find a positive number that is 20 less than its square
the equation for this statement:
x = x^2 - 20
0 = x^2 - x - 20; a quadratic equation that we can factor to
(x-5)(x+4) = 0
The positive solution only
x = 5